The angular position of a point on a rotating wheel is given by θ = 1.6 + 4.2t2 + 1.7t3, where θ is in radians and t is in seconds

(a) At t = 0, what is the point's angular position?

(b) At t = 0, what is the point's angular velocity?

(c) What is its angular velocity at t = 4.0 s?

(d) Calculate its angular acceleration at t = 2.0 s.

(e) Is its angular acceleration constant?
Answer #1

Given that 0 = 1.6+4.21? +1.743 - 0 a) nt to o=1.6+ 4.2x(0)2 + 1.7x1013 10=116 /oad And b) Angular velocity - do - d (1.6 +4.further if you have any doubt in any step comment below